OpenStudy (anonymous):

What is the x-intercept of -1/2x-5y=-10 ? Option A: 1/2 Option B: 2 Option C: 5 Option D: 20

OpenStudy (mathstudent55):

The x-intercept is where the graph crosses the x-axis. At that point, the y-coordinate is zero. Let y = 0 and solve for x.

OpenStudy (anonymous):

how do you solve

OpenStudy (mathstudent55):

Here's the equation: -1/2x-5y=-10 Now replace y with 0: \( -\dfrac{1}{2}x - 5y = 0\) \(-\dfrac{1}{2}x - 5 \cdot 0 = -10 \) \(-\dfrac{1}{2}x - 0 = -10 \) \(-\dfrac{1}{2}x = -10 \) Since x in beiung multiplied by -1/2, multiply both sides by -2 to find x.